# On the complex moment problem as a dynamic inverse problem for a discrete system

### v1

@@
Abstract. We consider the complex moment problem, that is the problem of constructing a positive Borel measure on $\mathbb{C}$ from a given set of moments. We relate this problem to the dynamic inverse problem for the discrete system associated with the complex Jacobi matrix. We show how the characterization of dynamic inverse data in solving the inverse problem provides sufficient conditions for solving the complex moment problem.
@@
In last section we describe connections between the IPs for (1.4) and (1.3) and complex moment problem and introduce a discrete Borel measure on $\mathbb{C}$, concentrated on the finite set of point, associated with (1.4), which is a solution to the truncated complex moment problem. Letting $N$ tend to infinity in (1.4) we obtain a sequence of measures that converges to a solution to the complex moment problem.
@@
(2.1) $UHU^{\top}=D=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$,
where $d_{i}\geqslant 0,$ $i=1\ldots,n$.
We apply this theorem for the matrix $A^{N}$: there exist unitary $U^{N}$ such that $U^{N}A^{N}\left(U^{N}\right)^{\top}$ as in the theorem. Note that in our case all $d_{i}>0$, since otherwise $A^{N}u=0$ implies $\det A^{N}=0$, which in turn implies linear dependence of some rows or columns of $A^{N}$, which is impossible. At the same time some of the $d_{i}$ can coincide. In the latter case we modify the unitary matrix as follows: assume for example that $d_{1}=d_{2}$, then, taking
@@
$U_{1}^{N}A^{N}\left(U_{1}^{N}\right)^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}e^{i\varphi}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix}$.
Using these arguments we can make all $d_{i}$ in the representation (2.1) distinct (but some of them can be complex). To summarize what we did so far, in what follows we choose unitary $U$ (we drop $N$) such that
(2.2) $UA^{N}U^{\top}=\begin{pmatrix}d_{1}&0&\ldots&0\\ 0&d_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&d_{n}\end{pmatrix},\quad d_{i}\in\mathbb{C}\backslash\{0\},d_{i}\not=d_{j},\,i\not=j,\,i,j=1\ldots,N.$
@@
(2.3) $A^{N}\hat{u}^{i}e_{i}=d_{i}\overline{\hat{u}^{i}},\quad A^{N}\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}=d_{i}\begin{pmatrix}\overline{\hat{u}^{i}_{1}}\\ \overline{{u}^{i}_{2}}\\ \ldots\\ \overline{\hat{u}^{i}_{n}}\end{pmatrix}$.
@@
so we formally add two values: $u^{i}_{0}=u^{i}_{N+1}=0$ and normalize vectors such that $u^{i}_{1}=1$.
@@
    Calculate parameters $d_{i}$, $\rho_{i}$, $H_{ki}$, $\omega_{i}$ using factorization (2.2) and formulas (2.7), (2.8), (2.12).

### v2

@@
Abstract. We consider the complex moment problem, that is the problem of constructing a positive Borel measure on $\mathbb{C}$ from a given set of moments. We relate this problem to the dynamic inverse problem for the discrete system associated with the complex Jacobi matrix. We propose a method that allows one to construct a discrete measure which is a solution to the truncated moment problem, we also show how the characterization of dynamic inverse data in solving the inverse problem provides sufficient conditions for solving the full complex moment problem.
@@
In last section we describe connections between the IPs for (1.4) and (1.3) and complex moment problem and introduce a discrete Borel measure on $\mathbb{C}$, concentrated on the finite set of point, associated with (1.4), which is a solution to the truncated complex moment problem. We emphasize that we propose a method that allows one to construct a discrete measure, i.e. find points of its support and determine the masses at these points. Letting $N$ tend to infinity in (1.4) we obtain a sequence of measures that converges to a solution to the full complex moment problem.
@@
(2.1) $UHU^{\top}=D=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix}$,
where $\hat{d}_{i}\geqslant 0,$ $i=1\ldots,n$.
We apply this theorem for the matrix $A^{N}$: there exist unitary $U^{N}$ such that $U^{N}A^{N}\left(U^{N}\right)^{\top}$ as in the theorem. Note that in our case all $\hat{d}_{i}>0$, since otherwise $A^{N}u=0$ implies $\det A^{N}=0$, which in turn implies linear dependence of some rows or columns of $A^{N}$, which is impossible. At the same time some of the $\hat{d}_{i}$ can coincide. In the latter case we modify the unitary matrix as follows: assume for example that $\hat{d}_{1}=\hat{d}_{2}$, then, taking
@@
$U_{1}^{N}A^{N}\left(U_{1}^{N}\right)^{\top}=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}e^{i\varphi}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix}$.
Using these arguments we can make all $\hat{d}_{i}$ in the representation (2.1) distinct (but some of them can be complex). To summarize what we did so far, in what follows we choose unitary $U$ (we drop $N$) such that
(2.2) $UA^{N}U^{\top}=\begin{pmatrix}\hat{d}_{1}&0&\ldots&0\\ 0&\hat{d}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&0&\ldots&\hat{d}_{n}\end{pmatrix},\quad\hat{d}_{i}\in\mathbb{C}\backslash\{0\},\hat{d}_{i}\not=\hat{d}_{j},\,i\not=j,\,i,j=1\ldots,N.$
@@
(2.3) $A^{N}\hat{u}^{i}=\hat{d}_{i}\overline{\hat{u}^{i}},\quad A^{N}\begin{pmatrix}\hat{u}^{i}_{1}\\ {\hat{u}}^{i}_{2}\\ \ldots\\ {\hat{u}}^{i}_{n}\end{pmatrix}=\hat{d}_{i}\begin{pmatrix}\overline{\hat{u}^{i}_{1}}\\ \overline{{u}^{i}_{2}}\\ \ldots\\ \overline{\hat{u}^{i}_{n}}\end{pmatrix}$.
@@
so we formally add two values: $u^{i}_{0}=u^{i}_{N+1}=0$ and normalize vectors such that $u^{i}_{1}=1$. In this case we have (see (2.3))
$A^{N}\begin{pmatrix}\frac{\hat{u}^{i}_{1}}{\hat{u}^{i}_{1}}\\ \frac{\hat{u}^{i}_{2}}{\hat{u}^{i}_{1}}\\ \ldots\\ \frac{\hat{u}^{i}_{n}}{\hat{u}^{i}_{1}}\end{pmatrix}={d}_{i}\begin{pmatrix}\frac{\overline{\hat{u}^{i}_{1}}}{\overline{\hat{u}^{i}_{1}}}\\ \frac{\overline{\hat{u}^{i}_{2}}}{\overline{\hat{u}^{i}_{1}}}\\ \ldots\\ \frac{\overline{\hat{u}^{i}_{n}}}{{\overline{\hat{u}^{i}_{1}}}}\end{pmatrix}$,
where we introduced the notation
(2.4) $d_{i}:=\hat{d_{i}}\frac{\overline{\hat{u}^{i}_{1}}}{{\hat{u}^{i}_{1}}},\quad i=1,\ldots,N.$
@@
    Calculate parameters $d_{i}$, $\rho_{i}$, $H_{ki}$, $\omega_{i}$ using factorization (2.2) and formulas (2.4), (2.8), (2.9), (2.13).